To give water and dinitrogen monoxide, 0.500g of ammonium nitrate decomposes. How much water is formed, in grams?
Answer #1

ANSWER : 0.225 gm of water is produced

EXPLANATION :

ammonium nitrate decomposes and produced water and dinitrogen monoxide(commonly known as laughing gas)

NH4NO3 ====> N2O + 2H2O

above reaction said that 1 mole of ammonium nitrate decomposes and produced 2 moles of water and 1 mole of dinitrogen monoxide.

molecular weight of ammonium nitrate = 80 gm/mole

mass of ammonium nitrate given = 0.500g

so mole of ammonium nitrate is;

mole=\frac{mass}{molecular weight}=\frac{0.5}{80}=0.00625

1 mole of ammonium nitrate produced 2 moles of water.

so 0.00625 moles of ammonium nitrate produced;

mole of H_{2}O=0.00625*2 =0.0125

MOLECULAR WEIGHT OF WATER = 18 gm/mole

so mass of water produced = mole* molecular weight = 0.0125 * 18 = 0.225 gm